# Ocean-E2E: Hybrid Physics-Based and Data-Driven Global Forecasting of Extreme Marine Heatwaves with End-to-End Neural Assimilation

### v2

@@ 3.3 End-to-end Neural Data Assimilation
Under realistic operational forecasting scenarios, the full initial ocean state $T_{0}$ (analysis field) cannot be directly observed. This necessitates a data assimilation (DA) process that optimally combines sparse observational data $T_{0}^{obs}$ with prior background estimates $T_{0}^{bg}$ from numerical models. Traditional DA methodologies require large computational resources, which imposes an inherent trade off between solution accuracy and operational feasibility boudier2023data .
@@
Our framework addresses these limitations by establishing a direct mapping from observational inputs $T_{0}^{obs}$ to assimilated states $T_{0}$ through deep learning. Instead of solving complex Bayesian inverse problems, we train neural networks to optimally combine sparse observations with background fields in one step. Specifically, we begin with the SSTA state at time $-t$ ($T_{-t}$), defining $[-t,0]$ as the assimilation window. We generate background fields ($T_{0}^{bg}$) by adding structured noise to $T_{-t}$:
$T_{0}^{bg}=\phi^{f}_{\theta}(T_{-t}+\varepsilon,A_{0:t})$ (21)
@@
where $\varepsilon$ denotes Perlin noises. The final analysis field is then produced by merging observations with background data:
$T_{0}=\phi^{a}_{\theta}(T_{0}^{obs},T_{0}^{bg})$ (22)
@@
Where $\phi^{a}_{\theta}$ is an edge-reparameterization and attention-guided network with Kirsch-guided reparameterization, which is capable of reconstructing SSTA fields from sparse observations liu2024real . The details of $\phi^{a}_{\theta}$ can be found in Appendix D.
4 Experiment
4.1 Benchmarks and Baselines
@@
We conduct the experiments on GLORYS12V1 reanalysis data lellouche2018copernicus , which provides daily mean data of sea surface temperature (SST) covering latitudes between -80° and 90° spanning between 1993 and 2021. The seasonal cycle has been removed from the original data in order to get the MHWs (SSTA) field shu2025advanced . The subset we use includes years from 1993 to 2021, which is 1993-2018 for training, 2019 for validing, and 2020-2021 for testing. The surface velocity of the ocean is obtain from satellie observation pujol2016duacs . The atmospheric variables are obtained from ECWMF Reanalysis v5 (ERA5) hersbach2020era5 . As shown in section 3.1, based on the physical governing process, we select three surface atmosphere variables to represent atmsopheric forcing $A_{0:t}$: 10m wind components ($U_{{\rm 10}}$, $V_{{\rm 10}}$) and 2m temperature ($T_{{\rm 2m}}$). More details, including observational data, can be found in Appendix B.
@@
Table 1: In global ocean simulation task, we compare the performance of our Ocean-E2E with 4 baselines. The average results for global SSTA of RMSE and CSI are recorded. A small RMSE ($\downarrow$) and a higher CSI ($\uparrow$) indicate better performance. The best results are in bold, and the second best are with underline. It should be noticed that some of the baseline models are not stable in this task and generate unrealisitic forecast results. As a result, RMSE exceeds 10 K is replaced with --.
Model Metric
@@
FourCastNet pathak2022FourCastNet 0.6836 0.2709 0.1755 0.1537 0.1423
CirT liu2025cirt 1.3496 0.0905 1.9249 0.0810 1.9932 0.0770 2.0321 0.0746
WenHai cui2025forecasting 0.5435 0.4202 0.7006 0.2805 0.7633 0.2447 0.8139 0.2209
ClimODE verma2024climode 0.7221 0.3263 0.8555 0.2091 0.8997 0.1896 0.1754
Ocean-E2E (ours) 0.5659 0.4285 0.6596 0.3230 0.6911 0.2874 0.7194 0.2580
Ocean-E2E (Promotion) 2.0% 5.8% 15.1% 9.5% 17.4% 11.6% 16.8%
Before introducing our experiments, there are two key concepts that should be clarified: ocean simulation and end-to-end ocean forecast. As pointed out in previous studiesshu2025advanced ; cui2025forecasting , the ocean simulation utilizes the current state of SSTA $T_{0}$ to forecast future state $T_{t}$ under the realistic atmosphere forcing ${\bf A}_{0:t}$. In contrast, the end-to-end forecast task includes assimilating the initial analysis field of SSTA $T_{0}$ and forecasting only based on the current atmosphere state ${\bf A}_{0}$.
We compare the simulation performance of Ocean-E2E with 3 data-driven models (Fourcastnetkurth2023fourcastnet , CirTliu2025cirt , and Wenhaicui2025forecasting ). Fourcastnet is an atmosphere backbone model with robust performance, CirT is a SOTA subseasonal-to-seasonal atmosphere prediction model, and Wenhai is the superior data-driven ocean model that outperforms traditional numerical models and data-driven models.
We also compare the end-to-end operational forecast ability against traditional numerical operational forecast system: ECWMF sub-seasonal to seasonal prediction (S2S). S2S is current superior subseasonal-to-seasonal weather forecast model that provides the forecast of atmosphere and ocean surface physics at a lead time of 40 days. The original resolution is 1.5° and we upsample the data to 1/2° for comparison.
4.2 Evalution Metric
We utilize two metrics, RMSE (Root Mean Square Error) and CSI (Critical Success Index), to evaluate the forecast performance. The RMSE represents the overall performance of our framework’s forecast, while CSI mainly focuses on the most extreme MHWs events. What’s more, we utilize assimilation bias (BIAS) to assess the systematic bias induced by our neural assimilation system. More details can be found in Appendix E.
@@ 4.2 Evalution Metric
Table 2: In operational global ocean forecast task, we compare the performance of our Ocean-E2E with S2S. The average results for global SSTA of RMSE are recorded. A small RMSE ($\downarrow$) indicates better performance. The best results are in bold, and the second best are with underline.
@@
Time Interval
10-day 20-day 30-day 40-day
S2S (2020) pathak2022FourCastNet 0.8140 0.8870 0.9514 0.9965
Ocean-E2E (2020) liu2025cirt 0.5750 0.6047 0.7514 0.8747
Ocean-E2E (Promotion) 29.4 % 31.8 % 21.02 % 12.2 %
S2S (2021) pathak2022FourCastNet 0.8261 0.8897 0.9450 0.9895
Ocean-E2E (2021) liu2025cirt 0.6414 0.7447 0.8147 0.8729
Ocean-E2E (Promotion) 22.4 % 16.3 % 13.7 % 11.7 %
@@
Figure 5: An illustration of the regional high-resolution simulation.
4.3 Evalution of ocean simulation performance
@@
4.4 Evaluation of the end-to-end neural assimilation and forecast performance
We conducted two sets of assimilation experiments using analysis fields from January 1, 2020 and January 1, 2021 as initial conditions, respectively. Both experiments employed a 6-day assimilation window to assimilate SSTA fields over a full year. As shown in Figure 4, the RMSE and BIAS evolution reveal that assimilation errors rapidly increase and stabilize during the assimilation process, demonstrating the stability of our assimilation framework. Notably, a persistent negative BIAS exists in our framework, indicating that our assimilation model systematically underestimates the intensity of marine heatwaves compared to GLORYS12V1 reanalysis data, as also evidenced in Figure 4a.
@@ 4.4 Evaluation of the end-to-end neural assimilation and forecast performance
4.5 Regional high-resolution simulation
As discussed in Section 3.1, our framework builds upon the GM90 theoretical framework, which is applicable to all eddy-permitting datasets (horizontal resolution: 10–100 km). To validate its performance, we implement a regional high-resolution simulation ($1/4^{\circ}$) of the Pacific Ocean. Figure 5 depicts the evolution of sea surface temperature anomaly (SSTA) during a 20-day simulation, governed by the equation:
$C_{t}-C_{0}=\int_{0}^{t}-\left(\mathbf{u}(t)+\mathbf{u}_{\theta}(\mathbf{u}(t),C)\nabla C+\dot{S}_{\theta}(\mathbf{A}(t),C)\right)dt,$ (23)
where $C_{t}$ and $C_{0}$ represent the SSTA at time $t$ and initial state, respectively. The results demonstrate that Ocean-E2E effectively resolves multi-scale physical processes: large-scale SSTA patterns align with air-sea interactions, while small-scale variations correspond to local advection effects quantified by $\int_{0}^{t}-\left(\mathbf{u}(t)+\mathbf{u}_{\theta}(\mathbf{u}(t),C)\right)\nabla C\,dt$. By explicitly modeling these components, the framework achieves rigorous reconstruction of $C_{t}-C_{0}$, confirming its capability to disentangle scale-dependent dynamics in MHWs bian2024scale .
@@ 4.5 Regional high-resolution simulation
Ocean-E2E w/o ADV 0.3884 0.1658
Ocean-E2E w/o S 0.3207 0.1504
Ocean-E2E 0.4285 0.2580
4.6 Ablation studies
To validate the effectiveness of the proposed framework, we perform an ablation analysis by sequentially removing two core components: the advection term $({\bf u}_{\theta}({\bf u},C)+{\bf u})\nabla C$ (denoted as ADV) and the source term $\dot{S}_{\theta}(\mathbf{A},C)$ (denoted as S). As demonstrated in Table 3, the Critical Success Index (CSI) exhibits a marked reduction when either component is omitted, confirming their essential roles in enhancing simulation accuracy, particularly for extreme marine heatwaves (MHWs). Notably, the source term contributes most significantly to forecasting performance, underscoring the critical influence of air-sea interaction dynamics on MHW evolution.
5 Conclusion
In this study, we present Ocean-E2E, a novel hybrid framework integrating data-driven and numerical approaches for end-to-end marine heatwave (MHW) forecasting. By explicitly resolving oceanic mesoscale advection and air-sea interactions through a differentiable dynamic kernel, the framework achieves stable 40-day global-to-regional MHW predictions. Notably, in 60-day global ocean simulations, Ocean-E2E demonstrated a significant improvement, reducing the RMSE by an average of 11.6% and increasing the CSI for extreme MHWs by an average of 9.2% compared to leading baseline models. Furthermore, in end-to-end operational forecast tasks, our framework outperformed the state-of-the-art S2S system, achieving an average RMSE reduction of approximately 10% over a 40-day forecast horizon using its own assimilated initial fields. Neural data assimilation further enables high-resolution forecasting independent of traditional numerical systems, while maintaining superior accuracy and stability, as evidenced by the rapid stabilization of assimilation errors. Experimental results comprehensively demonstrate its robustness across spatiotemporal scales, offering a promising tool for advancing MHW prediction and understanding climate extremes.
Acknowledgements
This work was supported by the National Natural Science Foundation of China (42125503, 42430602).
Appendix A Algorithm
@@
0:  Initial SSTA observation $T_{t}^{obs}$, analysis field $T_{t-1}^{a}$ (at time $t-1$), current state of atmospheric forcing $\mathbf{A}_{t}$ and ocean current velocity $\mathbf{O}_{t}$
0:  Current step SSTA analysis field $T_{t}^{a}$, Next step SSTA prediction $T_{t+1}$.
@@
4:     Perturb the initial SSTA field $T_{t-1}^{a}$ with $N=10$ Perlin noise $T_{t-1}^{a(i)}=T_{t-1}^{a}+\varepsilon^{(i)},i=1,2,...,N$.
5:     Feeding the perturbed ensembles into our trained forecast model to generate forecast ensemble: $T_{t}^{bg(i)}=\phi^{f}_{\theta}(T_{t-1}^{a(i)},\mathbf{A}_{t-1:t})$, which are also called background fields at time $t$.
6:     Using the neural assimilation network $\phi^{a}_{\theta}$ to produce the analysis field at time $t$: $T_{t}^{a}=\phi^{a}_{\theta}(T_{t}^{bg(i)},T_{t}^{obs})$
@@
8:     Calculate the future atmospheric forcing and ocean current velocity: $\mathbf{A}_{t:t+1}=\mathcal{M}_{\theta}(\mathbf{A}_{t}),\mathbf{O}_{t:t+1}=\mathcal{M}_{\theta}(\mathbf{O}_{t})$.
9:     Taking the current analysis field $T_{t}^{a}$, and integrating the equation (17) with $\mathbf{A}_{t:t+1}$ and $\mathbf{O}_{t:t+1}$ to get the forecast field $T_{t+1}$
@@
11:  return $Ocean-E2E$
Appendix B Data Details
B.1 Dataset
@@
Data used to train $\mathcal{M}_{\theta}$. As mentioned in Section 3.2, we need to pretrain an neural network $\mathcal{M}_{\theta}$ to model the evolution of the atmosphere forcing $\mathbf{A}$. In this study, we utilize one of the state-of-the-art weather forecast model: OneForecast [17]. The data used to train OneForecast is obtained from weatherbench2 [48] benchmark, which is a subset of ERA5 reanalysis data [24]. Details can be found in Table 4:
Table 4: The data used to train $\mathcal{M}_{\theta}$.
@@
Geopotential (Z) 13 1.5° 6h Global 1959$\sim$2021
Specific Humidity (Q) 13 1.5° 6h Global 1959$\sim$2021
Temperature (T) 13 1.5° 6h Global 1959$\sim$2020
U Component of Wind (U) 13 1.5° 6h Global 1959$\sim$2021
V Component of Wind (V) 13 1.5° 6h Global 1959$\sim$2021
10 Metre U Wind Component (U10M) 1 1.5° 6h Global 1959$\sim$2021
10 Metre V Wind Component (V10M) 1 1.5° 6h Global 1959$\sim$2021
2 Metre Temperature (T2M) 1 1.5° 6h Global 1959$\sim$2021
Mean Sea Level Pressure (MSLP) 1 1.5° 6h Global 1959$\sim$2021
Data used to train $\mathcal{N}_{\theta}$. As shown in Section 3.1, like the atmosphere forcing $\mathrm{A}$, we need another neural network $\mathcal{N}_{\theta}$ to model the ocean surface current dynamics. The data of surface current is obtained from satellie observation, which can be downloaded from https://data.marine.copernicus.eu/product/SEALEVEL_GLO_PHY_L4_MY_008_047. Details of this dataset can be found in Table 5:
Table 5: The data used to train $\mathcal{N}_{\theta}$.
@@
Eastward Surface Geostrophic Current Velocity ($\mathrm{U}_{g}$) 0.25° 24h Global 1993$\sim$2021
Northward Surface Geostrophic Current Velocity ($\mathrm{V}_{g}$) 0.25° 24h Global 1993$\sim$2021
Data used to train the hybrid data-driven network. The training of $\dot{\rm S}_{\theta}$ and $\textbf{u}_{\theta}$ involves three atmosphere variables and three ocean varibales, which can be found in Table 6. The SST dataset is selected from GLORYS12V1, which can be downloaded from https://data.marine.copernicus.eu/product/GLOBAL_MULTIYEAR_PHY_001_030.
Table 6: The data used to train $\dot{\rm S}_{\theta}$ and $\textbf{u}_{\theta}$.
@@
10 Metre U wind Component (U10M) 0.25° 24h Global 1993$\sim$2021
10 Metre V wind Component (V10M) 0.25° 24h Global 1993$\sim$2021
2 Metre Temperature (T2M) 0.25° 24h Global 1993$\sim$2021
Eastward Surface Geostrophic Current Velocity ($\mathrm{U}_{g}$) 0.25° 24h Global 1993$\sim$2021
Northward Surface Geostrophic Current Velocity ($\mathrm{V}_{g}$) 0.25° 24h Global 1993$\sim$2021
Sea Surface Temperature ($\mathrm{SST}$) 1/12° 24h Global 1993$\sim$2021
B.2 Data preprocessing
@@
For observational data, we utlilize the Hadley Centre Integrated Ocean Database (sparse observations) and Global Ocean ODYSSEA L4 Sea Surface Temperature (satellite observations), which can be downloaded from https://climatedataguide.ucar.edu/climate-data/hadiod-met-office-hadley-centre-integrated-ocean-database and https://data.marine.copernicus.eu/product/SST_GLO_PHY_L4_MY_010_044, respectively.
Appendix C Details of Dynamical Core of Ocean-E2E
@@
$C_{t}=\int_{0}^{t}-({\bf u}(t)+{\bf u}_{\theta}({\bf u}(t),C))\nabla C+\dot{\rm S}_{\theta}({\bf A}(t),C)dt+C_{0}$ (24)
@@
What’s more, unlike atmosphere variables, the ocean MHWs have coastline boundary, which should be treated properly to aviod any numerical instability. In our study, we mitigate this boundary effect through a boundart mask $\mathbf{M}$, where
$\mathbf{M}(i,j)=\begin{cases}1,&\text{if the grid locates on ocean}\\ 0,&\text{else}\end{cases}$ (27)

### v3

@@ Methods
As shown in the above section, our hybrid AI-physics model requires an initial condition $C_{0}$ to initialize our framework. However, under realistic operational forecasting scenarios, the full $C_{0}$ (analysis field) cannot be directly observed. This necessitates a data assimilation (DA) process that optimally combines sparse observational data $C_{0}^{obs}$ with prior background estimates $C_{0}^{bg}$ from numerical models, which requires large computational resources (Boudier et al. 2023).
@@
Our framework addresses these limitations by establishing a direct mapping from observational inputs $C_{0}^{obs}$ to assimilated states $C_{0}$ through deep learning. We utilized neural networks to optimally combine sparse observations with background fields in one step. Specifically, we begin with the SSTA state at time $\Delta t$ ($C_{-\Delta t}$), defining $\Delta t,0]$ as the assimilation window. We generate background fields ($C_{0}^{bg}$) by adding structured noise to $C_{-\Delta t}$:
$C_{0}^{bg}=\phi^{f}_{\theta}(C_{-\Delta t}+\varepsilon,\mathbf{A}_{0:\Delta t})$ (23)
@@
where $\varepsilon$ denotes Perlin noise. The final analysis field is then produced by merging observations with background data:
$C_{0}=\phi^{a}_{\theta}(C_{0}^{obs},C_{0}^{bg})$ (24)
@@
where $\phi^{a}_{\theta}$ is an edge-reparameterization and attention-guided network with Kirsch-guided reparameterization, which is capable of reconstructing SSTA fields from sparse observations (Liu et al. 2024). The details of $\phi^{a}_{\theta}$ can be found in Appendix D.
Once the initial analysis field $C_{0}$ is obtained, the model proceeds with forward assimilation over subsequent time steps to maintain accuracy in evolving forecasts. Starting from $C_{0}$, the framework iteratively advances the state by generating a background estimate for the next time step $C_{\Delta t}^{bg}$ using the forward operator and incorporating dynamical updates at a time interval $\Delta t$(i.e., assimilation window). For instance, at $i$-th time step, the background is computed as:
$C_{(i+1)\Delta t}^{bg}=\phi^{f}_{\theta}(C_{i\Delta t}+\varepsilon_{i},A_{i\Delta t:(i+1)\Delta t})$ (25)
followed by the assimilation step:
$C_{(i+1)\Delta t}=\phi^{a}_{\theta}(C_{(i+1)\Delta t}^{obs},C_{(i+1)\Delta t}^{bg})$ (26)
As we will demonstrate in the following paragraph, this autoregressive process can generate robust assimilation fields over hundreds of days.
Experiments
We designed comprehensive experiments to evaluate our model. Our evaluation is guided by the following key questions: RQ1: Overall Performance: Can Ocean-E2E outperform state-of-the-art models with higher consistency and better performance in normal and extreme MHW events? RQ2: End-to-end Forecasting: Can Ocean-E2E run independently of numerical model while preserving a high accuracy in real-world MHWs forecast? RQ3: High-resolution Multiscale Prediction: Can our framework achieve a satisfying performance in regional high-resolution simulations, where complicated oceanic multi-scale dynamics pose more challenges to the model’s forecast skill and physical consistency?
Benchmarks and Baselines
@@
We conduct the experiments on GLORYS12V1 reanalysis data (Lellouche et al. 2018), which provides daily mean data of sea surface temperature (SST) covering latitudes between -80° and 90° spanning between 1993 and 2021. The seasonal cycle has been removed from the original data in order to get the MHWs (SSTA) field (Shu et al. 2025). The subset we use includes years from 1993 to 2021, which is 1993-2018 for training, 2019 for validating, and 2020-2021 for testing. The surface velocity of the ocean is obtained from satellite observation (Pujol et al. 2016). The atmospheric variables are obtained from ECWMF Reanalysis v5 (ERA5) (Hersbach et al. 2020). As shown in section 3.1, based on the physical governing process, we select three surface atmosphere variables to represent atmsopheric forcing $A_{0:t}$: 10m wind components ($U_{{\rm 10}}$, $V_{{\rm 10}}$) and 2m temperature ($T_{{\rm 2m}}$). More details, including observational data, can be found in Appendix B.
@@
Table 1: In the global ocean simulation task, we compare the performance of our Ocean-E2E with various baselines. The average results for global SSTA (Sea Surface Temperature Anomaly) are measured using RMSE and CSI. Lower RMSE ($\downarrow$) and higher CSI ($\uparrow$) indicate better performance. The best results are in bold, and the second-best are underlined. Note that some baseline models are unstable and produce unrealistic forecasts, with RMSE exceeding 10 K replaced by —.
Model Category Metric
@@
Ocean Forecasting Models
[…]
Operator Learning Models
[…]
Computer Vision Backbones
[…]
Spatiotemporal Models
[…]
Promotion 2.0% 5.8% 15.1% 9.5% 17.4% 11.6% 16.8%
@@
Table 2: In operational global ocean forecast task, we compare the performance of our Ocean-E2E with S2S. The average results for global SSTA of RMSE are recorded for 2020 and 2021. A small RMSE ($\downarrow$) indicates better performance. The best results are in bold.
@@
2020 2021
10-day 20-day 30-day 40-day 10-day 20-day 30-day 40-day
S2S 0.8140 0.8870 0.9514 0.9965 0.8261 0.8897 0.9450 0.9895
Ocean-E2E (Ours) 0.5750 0.6047 0.7514 0.8747 0.6414 0.7447 0.8147 0.8729
Ocean-E2E (Promotion) 29.4% 31.8% 21.02% 12.2% 22.4% 16.3% 13.7% 11.7%
@@
Figure 5: Results of our regional high resolution MHWs simulation in western Atlantic Ocean. a)-c) Snapshots of our simulation, where yellow parts indicate MHWs. d)-e) RMSE and CSI of our Ocean-E2E and other baselines.
Before introducing our experiments, there are two key concepts that should be clarified: ocean simulation and end-to-end ocean forecast. As pointed out in previous studies(Shu et al. 2025; Cui et al. 2025), the ocean simulation utilizes the current state of SSTA $C_{0}$ to forecast future state $C_{t}$ under the realistic atmosphere forcing ${\bf A}_{0:t}$. In contrast, the end-to-end forecast task includes assimilating the initial analysis field of SSTA $T_{0}$ and forecasting only based on the current atmosphere state ${\bf A}_{0}$.
We compare the simulation performance of Ocean-E2E with 4 main catergories of data-driven models: ocean forecasting models (Fourcastnet (Kurth et al. 2023), CirT (Liu et al. 2025), Wenhai (Cui et al. 2025) and ClimODE (Verma, Heinonen, and Garg 2024)), operator learning models (FNO (Li et al. 2021), CNO (Raonic et al. 2023) and LSM (Wu et al. 2023)), computer vision backbones (U-Net (Ronneberger, Fischer, and Brox 2015), ResNet (He et al. 2016) and DiT (Peebles and Xie 2023)) and spatiotemporal models (ConvLSTM (Shi et al. 2015), SimVP (Tan et al. 2022) and PastNet (Wu et al. 2024b)).
We also compare the end-to-end operational forecast ability against traditional numerical operational forecast system: ECWMF sub-seasonal to seasonal prediction (S2S). S2S is currently the superior subseasonal-to-seasonal weather forecast model that provides the forecast of atmosphere and ocean surface physics at a lead time of 40 days. The original resolution is 1.5°, and we upsample the data to 1/2° for comparison.
Evalution Metric
We utilized two metrics, RMSE (Root Mean Square Error) and CSI (Critical Success Index), to evaluate the forecast performance. The RMSE represents the overall performance of our framework’s forecast, while CSI mainly focuses on the most extreme MHW events. What’s more, we utilize assimilation bias (BIAS) to assess the systematic bias induced by our neural assimilation system. More details can be found in Appendix E.
Evalution of ocean simulation performance (RQ1)
@@
Evaluation of the end-to-end neural assimilation and forecast performance (RQ2)
We conducted two sets of assimilation experiments using analysis fields from January 1, 2020 and January 1, 2021 as initial conditions, respectively. Both experiments employed a 6-day assimilation window to assimilate SSTA fields over a full year. As shown in Figure 4, the RMSE and BIAS evolution reveal that assimilation errors rapidly increase and stabilize during the assimilation process, demonstrating the stability of our assimilation framework.
@@
Regional high-resolution simulation (RQ3)
As discussed in Section 3.1, our framework builds upon the GM90 theoretical framework, which is also applicable to high-resolution data. To validate its performance, we implement a regional high-resolution simulation ($1/12^{\circ}$) of the western Atlantic Ocean. Figure 5 a)-c) illustrate a comparison between Ocean-E2E and baselines. Compared to the ground truth, our model has more physical details with more extreme MHW events (yellow parts). Evaluation based on RMSE and CSI (Figure 5 d)-e)) also confirmed that Ocean-E2E achieves a state-of-the-art performance in regional simulations.
@@
$\dot{S}_{\theta}$ (S only) 0.3884 0.1658
$(\mathbf{u}_{\theta}+\mathbf{u}_{g})\nabla C$ (ADV only) 0.3207 0.1504
$\mathbf{u}_{g}\nabla C$ (pure numerical)
$(\mathbf{u}_{\theta}+\mathbf{u}_{g})\nabla C+\dot{S}_{\theta}$ (Ours) 0.4285 0.2580
Ablation studies
To validate the effectiveness of the proposed framework, we perform an ablation analysis by sequentially removing core components in our global simulation task. Since our main contribution is the framework rather than neural network design, we choose the two key components of our framework: the advection term $({\bf u}_{\theta}({\bf u}_{g},C)+{\bf u}_{g})\nabla C$ (denoted as ADV) and the source term $\dot{S}_{\theta}(\mathbf{A},C)$ (denoted as S). As demonstrated in Table 3, the Critical Success Index (CSI) exhibits a marked reduction when either component is omitted, confirming their essential roles in enhancing simulation accuracy, particularly for extreme marine heatwaves (MHWs). Notably, the source term contributes most significantly to forecasting performance, underscoring the critical influence of air-sea interaction dynamics on MHW evolution. Moreover, if the neural networks are removed from the model (i.e., a pure numerical model), the model collapses quickly due to the numerical instability caused by large horizontal gradients in the advection terms. This stresses the importance of the GM ‘bolus’ velocity, which acts as a subgrid ‘damping’ term.
Conclusions
In this study, we introduce Ocean-E2E, a hybrid framework merging data-driven and numerical methods for end-to-end marine heatwave (MHW) forecasting. By resolving mesoscale advection and air-sea interactions with a dynamic kernel, it delivers stable 40-day global-to-regional MHW predictions in both simulation and forecast tasks. Neural assimilation enables independent high-resolution forecasting with enhanced accuracy and stability, via rapid error stabilization. Results confirm its robustness across scales, advancing MHW prediction and climate insights.
Appendix A Appendix A. Algorithm
@@ Conclusions
0: Initial SSTA observation $C_{t}^{obs}$, analysis field $C_{t-1}^{a}$ (at time t1t-1), current state of atmospheric forcing $\mathbf{A}_{t}$ and ocean current velocity $\mathbf{O}_{t}$ .
0: Current step SSTA analysis field $C_{t}^{a}$, Next step SSTA prediction $C_{t+1}$.
@@
4:  Perturb the initial SSTA field $C_{t-1}^{a}$ with $N=10$ Perlin noise $C_{t-1}^{a(i)}=C_{t-1}^{a}+\varepsilon^{(i)},i=1,2,...,N$.
5:  Feeding the perturbed ensembles into our trained forecast model to generate forecast ensemble: $C_{t}^{bg(i)}=\phi^{f}_{\theta}(C_{t-1}^{a(i)},\mathbf{A}_{t-1:t})$, which are also called background fields at time $t$.
6:  Using the neural assimilation network $\phi^{a}_{\theta}$ to produce the analysis field at time $t$: $C_{t}^{a}=\phi^{a}_{\theta}(C_{t}^{bg(i)},C_{t}^{obs})$
@@
8:  Calculate the future atmospheric forcing and ocean current velocity: $\mathbf{A}_{t+1}=\mathcal{N}_{\theta}(\mathbf{A}_{t}),\mathbf{O}_{t+1}=\mathcal{M}_{\theta}(\mathbf{O}_{t})$.
9:  Taking the current analysis field $C_{t}^{a}$, and integrating the equation (17) with $\mathbf{A}_{t:t+1}$ and $\mathbf{O}_{t:t+1}$ to get the forecast field $C_{t+1}$
@@
11:return OceanE2EOcean-E2E
Appendix B Appendix B. Data Details
Dataset
@@
Data used to train $\mathcal{N}_{\theta}$. As mentioned in Section 3.2, we need to pretrain an neural network $\mathcal{N}_{\theta}$ to model the evolution of the atmosphere forcing $\mathbf{A}$. In this study, we utilize one of the state-of-the-art weather forecast model: OneForecast (Gao et al. 2025). The data used to train OneForecast is obtained from weatherbench2 (Rasp et al. 2024) benchmark, which is a subset of ERA5 reanalysis data (Hersbach et al. 2020). Details can be found in Table 4.
Table 4: The data used to train $\mathcal{N}_{\theta}$.
@@
[…]
Data used to train $\mathcal{M}_{\theta}$. As shown in Section 3.1, like the atmosphere forcing $\mathbf{A}$, we need another neural network $\mathcal{M}_{\theta}$ to model the ocean surface current dynamics. The data of surface current is obtained from satellie observation, which can be downloaded from https://data.marine.copernicus.eu/product/SEALEVEL˙GLO˙PHY˙L4˙MY˙008˙047. Details of this dataset can be found in Table 5.
Table 5: The data used to train $\mathcal{M}_{\theta}$.
@@
[…]
Data used to train the hybrid data-driven network. The training of $\dot{\rm S}_{\theta}$ and $\textbf{u}_{\theta}$ involves three atmosphere variables and three ocean varibales, which can be found in Table 6. The SST dataset is selected from GLORYS12V1, which can be downloaded from https://data.marine.copernicus.eu/product/GLOBAL˙MULTIYEAR˙PHY˙001˙030. For regional high-resolution simulation, we choose a regional subset of GLORYS12V1 (Table 7).
Table 6: The data used to train $\dot{\rm S}_{\theta}$ and $\textbf{u}_{\theta}$ (global simulation).
@@
[…]
Table 7: The data used to train $\dot{\rm S}_{\theta}$ and $\textbf{u}_{\theta}$ (regional simulation).
Variable
Name
Spatial
Resolution
Dt
Lat-Lon
Range
Time
[…]
Observational data. For observational data, we utlilize the Hadley Centre Integrated Ocean Database (sparse observations) and Global Ocean ODYSSEA L4 Sea Surface Temperature (satellite observations), which can be downloaded from https://climatedataguide.ucar.edu/climate-data/hadiod-met-office-hadley-centre-integrated-ocean-database and https://data.marine.copernicus.eu/product/SST˙GLO˙PHY˙L4˙MY˙010˙044, respectively.
Data preprocessing
@@
Appendix C Appendix C. Details of Dynamical Core of Ocean-E2E
@@
$C_{t}=\int_{0}^{t}-({\bf u}_{g}(t)+{\bf u}_{\theta}({\bf u}_{g}(t),C))\nabla C+\dot{\rm S}_{\theta}({\bf A}(t),C)dt+C_{0}$ (27)
@@
What’s more, unlike atmosphere variables, the ocean MHWs have coastline boundary, which should be treated properly to aviod any numerical instability. In our study, we mitigate this boundary effect through a boundary mask $\mathbf{M}$, where
[…]